OpenStudy (anonymous):

(x^2 + 3x - 10) / (6x^2 - 24x) multiplied by (2x^2 - 4x) / (x + 5)

OpenStudy (anonymous):

\[(x ^{2}+3x-10)/(6x ^{2}-24x)*(2x ^{2}-4x)/(x+5)\]

OpenStudy (anonymous):

Is that what it looks like?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i worked out an answer but want to make sure i understand how to do these problems...

OpenStudy (anonymous):

okay do you want to show me?

OpenStudy (anonymous):

yes, thanks :)

OpenStudy (anonymous):

Okay:)

OpenStudy (anonymous):

first step got (x + 5) (x - 2) / 6x (x - 4) multiplied by 2x (x - 2) / (x + 5)

OpenStudy (anonymous):

answer (x - 2) (x - 2) / 3 (x - 4)

OpenStudy (anonymous):

So your just simplifying the equation?

OpenStudy (anonymous):

I think so, the directions just say to complete the problems.

OpenStudy (anonymous):

It looks fine to me, Do you think the x-2's can be simplified more?

OpenStudy (anonymous):

I wasn't sure.

OpenStudy (anonymous):

Okay than I wouldn't worry about it :) Congrats.

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

would you help with another one?

OpenStudy (anonymous):

BR goes to you:) I could spare some more time :P

OpenStudy (anonymous):

thanks :) factor 27y^3 - 8x^6

OpenStudy (anonymous):

(3y)^3 - 2^3 (x^2)^3 ... but can this be factored as a cube?

OpenStudy (anonymous):

a3 + b3 = (a + b)(a2 – ab + b2) a3 – b3 = (a – b)(a2 + ab + b2)

OpenStudy (anonymous):

It seems like it should but when I substitute and then check it, it doesn't work out so not sure where I am making a mistake.

OpenStudy (anonymous):

Sorry I meant: a^3+b^3=(a+b)(a^2-ab+b^2) a^3-b^3=(a-b)(a^2+ab+b^2) I don't believe it can be factored as a cube.

OpenStudy (anonymous):

oh, ok thanks :)

OpenStudy (anonymous):

No Problem, Hoped I helped :)

OpenStudy (anonymous):

You did :)